At the end of April, Cavy Company had completed Jobs 766 and 765. According to the individual job cost sheets the information is as follows:

Job Direct Materials Direct Labor Machine Hours

Job 765 $10,080 $3,024 36

Job 766 $14,877 $4,698 87

Job 765 consisted of 216 units, and Job 766 consisted of 261 units.

Assuming that the predetermined overhead rate is applied by using machine hours at a rate of $137 per hour.

a. Determine the balance on the job cost sheets for each job.

Job 765 $ ______

Job 766 $ ______

b. Determine the cost per unit at the end of April. Round to the nearest cent.

Job 765 $ ______

Job 766 $ ______
Answer #1

a) Job Cost sheet

Job 765 Job 766
Direct material 10080 14877
Direct labor 3024 4698
Overhead 36*137 = 4932 87*137 = 11919
Total 18036 31494

b) Cost per unit

Job 765 Job 766
Direct material 10080 14877
Direct labor 3024 4698
Overhead 36*137 = 4932 87*137 = 11919
Total 18036 31494
Unit 216 261
Unit cost 83.50 120.67